OpenStudy (anonymous):

-2/sqr2 divided by 1/4sqr2 How to solve it?

OpenStudy (agent0smith):

\[\huge \frac{ -2 }{ \sqrt2 } \div \frac{ 1 }{ 4\sqrt2 }\] Dividing by a fraction is equivalent to multiplying by the inverse of the fraction, so this becomes: \[\huge \frac{ -2 }{ \sqrt2 } \times \frac{ 4\sqrt2 }{ 1 }\]

OpenStudy (anonymous):

could you solve it further please ?

OpenStudy (agent0smith):

All you need to do now is multiply the numerators together, and the denominators together.

OpenStudy (anonymous):

so the top will be -8 root 2 divided by root 2 ?

OpenStudy (agent0smith):

Correct. Now you can simplify.

OpenStudy (anonymous):

will the answer be -8?

OpenStudy (agent0smith):

Yep.

OpenStudy (anonymous):

but my calculator gives me -4 and the answer is -4

OpenStudy (agent0smith):

http://www.wolframalpha.com/input/?i=-2%2F%28sqr2%29+divided+by+%281%2F%284sqr2%29%29 Unless you've written the problem incorrectly, the answer is -8.

OpenStudy (anonymous):

thanks anyway.

OpenStudy (anonymous):

Thats why I was confused. I did it by hand the checked with my cal and I got different answers .

OpenStudy (agent0smith):

\[\huge \frac{ -2 }{ \sqrt2 } \div \frac{ 1 }{ 4\sqrt2 }\] \[\huge \frac{ \frac{ -2 }{ \sqrt2 }} { \frac{ 1 }{ 4\sqrt2 } }\] As long as the problem looks like either of these, it's equal to -8.